Fermi National Laboratory


Accelerator Update - July 11-July 14

Friday July 11
The day shift began with operators preparing the Booster and the Main Injector (MI) for an access, and with Operations monitoring store 2774.

By a little after 8 AM technicians had begun work on a MI kicker and power supply (MH2), and a Linac RF station (LRF2).

Operations terminated store 2774 at 1:04 PM. Operators prepared the Collider Detector Facility (CDF) and D-zero (D0) experimental collision halls for an access. Both experiments were out of access by 2:30 PM.

Operations put the Tevatron (TeV) through a dry squeeze at 3:47 PM.

Operations established beam to the Linac dump at 3:58 PM.

The Antiproton Source (Pbar) began stacking at 5:14 PM.

Operations resumed sending beam to MiniBooNE at 5:33 PM.

Operations put the TeV into shot setup at 6:25 PM.

The Booster suffered a vacuum burst at 7:47 PM. Everything was reset within five minutes.

At 8:41 PM, Operations established store 2780 with an initial luminosity of 37.90E30. Pbar resumed stacking.

A MI expert began conducting damper studies at 9 PM.

Saturday July 12
The midnight shift began with Operation monitoring store 2780, with Pbar stacking, and with MiniBooNE taking beam.

The MI damper studies were finished at 1 AM.

Store 2780 luminosity took a step down at 1:40 AM due to separator sparking.

At 1:52 AM, operators discovered a Tevatron (TeV) RF station (TRF5) not working. Operators called an expert who suggested that the station's water heater be turned off.

The I- Source started sparking a little at 3:40 AM.

Two Switchyard experts tried to conduct studies at 9:53 AM, but many problems delayed them.

The I- Source tripped off at 10:38 AM. It reset okay.

Stacking stopped at 11:49 AM due to a kicker problem (A:IKIK). Operators called in an expert who, at 1 PM, found a bad connector.

A MI expert began conducting studies at 12:53 PM. She ended her studies four hours later.

A Booster expert arrived at 2 PM to investigate a problem with a kicker (MKS04). It took about two hours to complete repairs.

A Linac RF station tripped off at 3:14 PM. Operators had the station back on by 3:41 PM, but it tripped off again just a few minutes later. Operations rebooted several systems.

At 4:38 PM, a TeV RF station (TRF4) tripped off and terminated store 2780. Fortunately it was a clean abort with no quench.

Operations put the TeV through a dry squeeze at 5:09 PM.

A TeV RF expert arrived to work on TRF5 at 5:42 PM. He replaced a power supply, among other things, and had the station working by 6:30 PM.

Operations adjusted the Booster coalescing at 6 PM. Operators noticed another problem and contacted a Booster expert to come in. He said the problem was with a kicker (MKS03) and that it wouldn't last long. The repair, which required a tunnel access, was scheduled for 8:30 AM, Sunday morning.

Operations put the TeV into shot setup at 8:36 PM. At 11 PM, while performing a reverse proton tune up, problems occurred. Operation consulted with experts for advice.

A Linac RF station (LRF4) tripped off at 11:22 PM for the first of three times within fifteen minutes.

Sunday July 13
The midnight shift began with Operations adjusting TeV chromaticity, and with MiniBooNE taking beam.

A Linac RF station tripped off at 12:09 AM. It reset okay.

Operation halted shot setup at 1 AM because the TeV problems were not understood.

A Booster power supply (ORBMP) tripped off at 1:46 AM due to an interlock fault. It tripped off three more times within the next forty-five minutes. Operators adjusted a limit on a LCW flow switch.

Operations resumed shot setup at 2:27 AM.

Operations established store 2783 at 3:40 AM. Its initial luminosity was 32.62E30. Pbar resumed stacking.

Operators rebooted the Switchyard front end at 7:22 AM to restore a link, but it didn't work. Operators contacted a controls expert who swapped out some equipment, but there were still problems. Eventually, techs found a bad crate power supply and replaced it at 10:52 AM.

Operations turned off the Booster at 7:32 AM in preparation for the tunnel access and repair of kicker MKS03. This halted all stacking and MiniBooNE beam. The repairs began at 8:44 AM. At 9:18 AM, an expert stated that the kicker was beyond repair and had to be replaced.

The Booster work was completed by 11 AM. Stacking resumed at 11:13 AM, and beam to MiniBooNE at 11:27 AM.

The Booster suffered a vacuum burst at 8:06 PM. It reset okay.

At 9:49 PM, the Booster ORBMP power supply tripped off again on an interlock fault. It tripped off two more times within the next thirty minutes. At 10:33 PM, operators adjusted two LCW flow switches.

The TeV store 2783 aborted at 10:16 PM. A number of TeV RF stations tripped off. TRF4 sparked and may have caused the other station to also spark. There was no quench. An expert came in and began investigating the TeV damper system.

Pbar resumed stacking at 10:33 PM.

Booster RF stations 13 and 14 tripped off at 11:45 PM on a high vacuum fault. The fault would not clear.

Monday July 14
The midnight shift began with a Booster expert investigating the RF stations 13 and 14, and with a TeV expert investigating the RF problem.

MiniBooNE beam returned at 12:25 AM.

At 12:30 AM, a TeV RF expert replaced a bad power supply.

Operations put the TeV through a squeeze with beam at 1:35 AM. At 2:13 AM, Operations put the TeV into shot setup.

Operations established store 2786 at 3:32 AM with an initial luminosity of 30.92E30. Pbar resumed stacking.

Today's Plan
The plan for today is to stack and store.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 7/15/2003   email Fermilab